Grade 11Physical Chemistry

Find the molecular formula of sulphur vapour when 3.2g of Sulphur is vapourized undeer conditions 450 degree C and 723mm pressure , and the vapour occupies a volume of 780mL.

Ipshita Prosad
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
The Reaction would be:
Total moles = moles of sulphur + moles of Suphur vapor.
                 = 3.2/32 + n
From ideal gas eqn. we get,
(723/760)*(780/1000) = (n+0.1)*0.0821*723
(n+0.1) = [(723/760)*(780/1000)]/[(0.0821*723)] = 0.0125
as n is coming negative and hence the moles can’t be nagtive the Reaction is not feasible here.